The following data from the 12/31/20xx financial statements of Designer Style Company:

Administrative Expenses           $   70,000        Interest Expense                                    $   5,000

Cost of Goods Sold                     500,000        Income Tax Expense                               42,000

Other Income                                 10,000         Net Sales                                             900,000

Total Assets (Beg bal.)                 900,000         Selling Expenses                                     100,000

Total Assets (End bal.)              1,100,000

What was the Gross Profit?

What was the Gross Profit Percentage?

What was the Total Assets Turnover?

What was the Income from Operations?

What was the Income before Income Taxes?

What was the Net Profit Percentage?
Answer #1

Gross profit:

Net sales 900,000
Less: Cost of goods sold -500,000
Gross profit 400,000

Gross profit percentage:

Gross profit percentage = (Gross profit ÷ Net sales ) x 100

= (400,000/900,000) x 100

= 44.44%

Total assets Turnover:

Average total assets = (Beginning balance + Ending balance)/2 = (900,000+1,100,00)/2 = 1,000,000

Total Assets turnover = Net sales ÷ Average total assets

= 900,000/1,000,000

= 0.90

Operating income:

Gross profit 400,000
Less: Operating expenses
Selling Expenses -100,000
Operating income 300,000

Income before income taxes:

Operating income 300,000
Non-operating income(expenses):
Other income 10,000
Interest expense -5,000
Income before income taxes 305,000

Net profit percentage:

Income before income taxes 305,000
Less: Income tax expense -42,000
Net profit 263,000

Net profit percentage = (Net profit ÷ Net sales) x 100

= (263,000/900,000) x 100

= 29.22%